OpenStudy (anonymous):

Find all solutions of the equation sec2x−2=0. The answer is A+Bkπ where k is any integer and 0

OpenStudy (anonymous):

Find all solutions of the equation sec2x−2=0. The answer is A+Bkπ where k is any integer and 0<A<π/2 A=? B=?

OpenStudy (anonymous):

sec2x = 2 so cos2x = 1/2 so 2x = 2*pi*k +- pi/3 so x = pi*k +- pi/6 since we require that 0 < A < pi/2, then we have to choose the + option so x = pi*k + pi/6 A = pi/6, B = 1

OpenStudy (anonymous):

ans doesnt seem to work for me

OpenStudy (phi):

I always change sec to 1/cos \[ \sec^2x−2=0 \] becomes \[ \frac{1}{\cos^2 x} -2 =0 \] add 2 to both sides \[ \frac{1}{\cos^2 x} -2 +2=0 +2\] simplify \[ \frac{1}{\cos^2 x} = 2\] "flip" both sides \[ \cos^2 x = \frac{1}{2} \] take the square root of both sides \[ \cos x = \sqrt{\frac{1}{2} } = \frac{1}{\sqrt{2}}= \frac{\sqrt{2}}{2}\]

OpenStudy (phi):

we should write that last line as \[ \cos x = \pm \frac{\sqrt{2}}{2} \] this is a well-known result.